40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 25 15:43:29 2016, ericq, Update, SUS, OSEM issues SRM_OSEMS.pngPRM_LR.png
    Reply  Thu Aug 25 20:07:35 2016, gautam, Update, SUS, OSEM issues - maybe resolved? 
    Reply  Tue Sep 6 18:21:13 2016, rana, Update, SUS, OSEM issues PRMswing.png
       Reply  Tue Sep 6 20:52:42 2016, gautam, Update, SUS, OSEM issues P9060254.JPG
          Reply  Wed Sep 7 17:23:26 2016, gautam, Update, SUS, PRM LR fixed for now IMG_3208.JPGIMG_3211.JPG
             Reply  Tue Oct 11 00:43:50 2016, ericq, Update, SUS, PRM LR problematic again PRM_LR_Problem.pngsince_fix.pnglast10.png
                Reply  Tue Oct 11 10:15:04 2016, Steve, Update, SUS, PRM LR problematic again PRMalfoiled.png
                   Reply  Tue Oct 11 13:30:49 2016, gautam, Update, SUS, PRM LR problematic again PRM_LR_decay.png
                      Reply  Wed Oct 12 13:34:28 2016, gautam, Update, SUS, PRM LR problematic again 08.png
Message ID: 12551     Entry time: Tue Oct 11 13:30:49 2016     In reply to: 12549     Reply to this: 12552
Author: gautam 
Type: Update 
Category: SUS 
Subject: PRM LR problematic again 

Perhaps the problem is electrical? The attached plot shows a downward trend for the LR sensor output over the past 20 days that is not visible in any of the other 4 sensor signals. The Al foil was shorting the electrical contacts for nearly 2 months, so perhaps some part of the driver circuit needs to be replaced? If so a Satellite Box swap should tell us more, I will switch the PRM and SRM satellite boxes. It could also be a dying LED on the OSEM itself I suppose. If we are accessing the chamber, we should come up with a more robust insulating cap solution for the OSEMs rather than this hacky Al foil + kapton arrangement.

The PRM and SRM Satellite boxes have been switched for the time being. I had to adjust some of the damping loop gains for both PRM and SRM and also the PRM input matrix to achieve stable damping as the PRM Satellite box has a Side sensor which reads out 0-10V as opposed to the 0-2V that is usually the case. Furthermore, the output of the LR sensor going into the input matrix has been turned off.

 

ELOG V3.1.3-